On October 1, 2021, the Allegheny Corporation purchased equipment for $314,000. The estimated service life of the equipment is 10 years and the estimated residual value is $6,000. The equipment is expected to produce 550,000 units during its life.

Required:
Calculate depreciation for 2021 and 2022 using each of the following methods. Partial-year depreciation is calculated based on the number of months the asset is in service.

3. Units of production (units produced in 2021, 28,000; units produced in 2022, 43,000). (Round "Depreciation per unit rate" answers to 2 decimal places.)
Answer #1

Depreciation expense per unit = (314,000-6,000)/550,000

= 0.56 per unit

Depreciation for

2021 = 28,000*0.56 = 15,680

2022 = 43,000*0.56 = 24,080